OpenStudy (kayders1997):

find the critical numbers of the function g(x)=x^3/4-2t^1/4

OpenStudy (tkhunny):

Is it really t^(1/4)? I'm questioning the 't'.

OpenStudy (kayders1997):

my bad it was x

OpenStudy (kayders1997):

I was looking at my book, it was originally t's but I don't like them so I change them to x

OpenStudy (tkhunny):

Not a great plan, really. Try putting tails on them and your handwriting should be more distinguishable.

OpenStudy (tkhunny):

Well, you'll need a first derivative, no?

OpenStudy (kayders1997):

right I got 3/4t^-1/4-1/2t^-3/4

OpenStudy (kayders1997):

for the first number I got t=0 if you want to go by t's

OpenStudy (tkhunny):

That's fine. Can you pretty it up a bit?

OpenStudy (kayders1997):

do you mean by factoring?

OpenStudy (kayders1997):

cause of thats what you mean you can take out a 1/4^-3/4(3t^1/2-2)

OpenStudy (tkhunny):

I don't know. \(\dfrac{3}{4t^{1/4}} - \dfrac{2}{t^{3/4}} = \dfrac{3t^{1/2} - 8}{4t^{3/4}}\) Kinda messy any way you look at it, but the solution is easier with only one denominator.

OpenStudy (kayders1997):

I don't understand where you got ohhhh just kidding

OpenStudy (kayders1997):

so now you have to set that equal to 0

OpenStudy (kayders1997):

I got it thank you for the help

OpenStudy (tkhunny):

How may solutions did you find?

OpenStudy (kayders1997):

2

OpenStudy (kayders1997):

t=0 and t=4/9

OpenStudy (tkhunny):

Are you sure it's not symmetric about the y-axis and there isn't another for negative t?

OpenStudy (kayders1997):

its right I looked in the back of the book

OpenStudy (tkhunny):

Books can be wrong. Good luck.

OpenStudy (kayders1997):

I'm pretty sure it's right, I mean thats what I got than checked the back